When heated, KClO3 decomposes into KCl and O2.

2KCLO3 ----> 2KCL +3O2

If this reaction produced 51.4 g of KCl, how much O2 was produced (in grams)?
1
Expert's answer
2018-04-07T08:18:11-0400
2KClO3 ----> 2KCl +3O2
According to the equation 2 moles of KCl and 3 moles of O2 are produced after the decomposition of 2 moles of KClO3.
Find how many moles of KCl were produced:
ν = m/M = 51.4/74.5 = 0.69 (mole);
Find how many moles of O2 were produced:
ν = 0.69 × 1.5 = 1.035 (mole) ;
Find the mass of O2 :
m = ν × M = 1.035 × 32 = 33.12 (g)
Answer
33.12 g of O2 were produced.